3 identical small balls ,each of mass 0.1 gm are suspended at one point on silk threads having a length of 20cm. what charges should be imparted to the balls for each thread to form an angle of 30° with the vertical?

Grade:12

1 Answers

Sher Mohammad IIT Delhi
9 years ago
vertical foce = mg
horizontal force = qe
tan (angle)= mg/qe
tan(30)=mg/qe
q= mg/tan(30)e
sher mohammad
faculty askiitians
b.tech, iit delhi